# Small-time asymptotics and the emergence of complex singularities for the KdV equation

### v1

@@ 2.2 Exponential asymptotics argument for limit $t\rightarrow 0^{+}$
$\times\cos\left(\frac{2}{3(3t)^{1/2}}(x^{2}+y_{0}^{2})^{3/4}\sin{\textstyle\frac{3}{2}}\phi+{\textstyle\frac{1}{4}}\phi\right),$ (18)
@@
these waves take the form (5). Thus, we see dispersive waves whose amplitude is exponentially small in both limits $x\rightarrow$-$\infty$ and $t\rightarrow 0^{+}$ (compared to the initial condition $u_{0}(x)$, which is assumed to be $\mathcal{O}(1)$), scaling as
@@
$x_{m}\sim-3\pi^{2/3}m^{2/3}t^{1/3},\quad m\rightarrow\infty,\quad t\rightarrow 0^{+},$ (19)
so that the speed of the waves decreases algebraically as $\dot{x}_{m}\sim-\pi^{2/3}(t/m)^{-2/3}$ and the wavelength increases as $x_{m+1}+x_{m}\sim 2\pi^{2/3}(t/m)^{1/3}$ as $t$ increases from zero. These scalings explain why it is so difficult to compute numerical solutions of the KdV model (1) accurately for small time using elementary techniques (such as finite differences), given the challenge of capturing fast-moving, small-amplitude waves that would immediately reflect off the boundary of any truncated domain. Such issues can be resolved, for example, via an inverse-scattering formulation, but such an approach is highly nontrivial [81] (and, of course, limited to integrable systems); we instead use the spin (stiff pde integrator) command [66] in Chebfun [35], which is good enough for our real-valued numerical solutions.
@@ 2.3 Numerical tests for the exponential asymptotics
To test (18) against numerics, we consider the small-time expansion
@@
i.e., we retain only the first two terms of the algebraic expansion (7) but include the leading-order exponential correction (18) (in order to capture the dominant oscillatory contribution). For the initial condition (4), we have
@@
and $y_{0}=1$. We also make use of a scaled version of (18), namely $U_{\mbox{\footnotesize scaled}}=U_{\mbox{\footnotesize dis}}/K$, where
$K=t^{-3/4}(x^{2}+y_{0}^{2})^{1/8}\mathrm{e}^{2(x^{2}+y_{0}^{2})^{3/4}\cos(3\phi/2)/(3(3t)^{1/2})}$ (20)
@@
For example, in figure 2(a), we show a numerical solution of (1), $u_{\mathrm{num}}$, computed with the $\mathrm{sech}^{2}$-type initial condition (4), with $A_{0}=-1/4$, for a representative early time, $t=0.02$. The very small dispersive waves we see in the inset on the left panel are propagating to the left. In the middle panel, we plot $u_{\mathrm{num}}-(u_{0}+tu_{1})$ at the specific time $t=0.02$, and compare with $U_{\mbox{\footnotesize dis}}$ from (18). We see for these parameter values, the comparison is very good, which is a strong test for both our numerics and asymptotics. As another test, we plot in the right panel the scaled versions $u_{\mathrm{num}}-(u_{0}+tu_{1})/K$ and $U_{\mbox{\footnotesize scaled}}=U_{\mbox{\footnotesize dis}}/K$, where $K$ is defined in (20). The agreement is excellent.
@@
Figure 2: (a) [left panel] Numerical solution of the KdV model (1) on the real line, computed at $t=0.02$ using the $\mathrm{sech}^{2}$-type initial condition (4) with $A_{0}=-1/4$; [middle panel] plots of $u_{\mathrm{num}}-(u_{0}+tu_{1})$ (blue) and $U_{\mbox{\footnotesize dis}}$ (red circles) versus $x$ using same parameters as in left panel; [right panel] plots of $\left(u_{\mathrm{num}}-(u_{0}+tu_{1})\right)/K$ (blue) and $U_{\mbox{\footnotesize scaled}}$, again with same parameters as left panel. (b) Same as (a), except that $A_{0}=-3/4$.
We have generated a number of other examples like that presented in figure 2(a) using the initial condition (4), computed for other values of $A_{0}<0$, and the agreement between numerics and asymptotics is also excellent. One such example is for $A_{0}=-3/4$, is shown in figure 2(b). This sweep of parameters includes the special cases $A_{0}=-N(N+1)$, where $N\in\mathbb{N}$, for which we have exact $N$-soliton solutions with no dispersive waves at all. For those special choices, the amplitude of the waves is predicted by (18) to vanish, since
$\cos\left(\frac{\pi}{2}\sqrt{1-4A_{0}}\right)=\cos\left(\frac{\pi}{2}\sqrt{(2N+1)^{2}}\right)=\cos\left(N\pi+\frac{\pi}{2}\right)=0$
for all $N\in\mathbb{N}$. Putting it together, we have presented in figure 3(a) a comparison of our numerical estimate of the amplitude of the wavelike term $\left(u_{\mathrm{num}}-(u_{0}+tu_{1})\right)/K$ (blue dots) with the asymptotic prediction $(2/3^{3/4}\pi^{1/2})\cos\left(\frac{\pi}{2}\sqrt{1-4A_{0}}\right)$ (red solid) for simulations that take the $\mathrm{sech}^{2}$-type initial condition (4). Given the very good agreement in this plot over a range of values of $A_{0}$, we are confident the approximation (18) is correctly describing the dispersive waves, at least for the initial condition (4), in the small-time limit.
@@
Figure 3: A numerically computed amplitude of the scaled dispersive waves $\left(u_{\mathrm{num}}-(u_{0}+tu_{1})\right)/K$ (blue dots) plotted for various values of the parameter $A_{0}$ at a fixed time $t=0.03$, compared with the asymptotic prediction $(2/3^{3/2}\pi^{1/2})\cos\left(\frac{\pi}{2}\sqrt{1-4A_{0}}\right)$ (red solid). (a) the $\mathrm{sech}^{2}$-type initial condition (4); (b) the initial condition (6); (c) the refined initial condition (22).
@@
Figure 4: (a) [left panel] Numerical solution of the KdV model (1) on the real line, computed at $t=0.02$ using the generic initial condition (6) with $A_{0}=-1/4$; [middle panel] plots of $u_{\mathrm{num}}-(u_{0}+tu_{1})$ (blue) and $U_{\mbox{\footnotesize dis}}$ (red circles) versus $x$ using same parameters as in left panel; [right panel] plots of $\left(u_{\mathrm{num}}-(u_{0}+tu_{1})\right)/K$ (blue) and $U_{\mbox{\footnotesize scaled}}$, again with same parameters as left panel. (b) Same as (a), except that $A_{0}=-3/4$. (c) Same as (a), except that $A_{0}=-2$.
@@
These types of numerical tests turn out to be slightly more subtle for the initial condition (6). For example, in figure 4 we show plots that are analogous to figure 2, except that in figure 4 we use the initial condition (6). Broadly speaking, we see very good agreement between numerics and asymptotics for $A_{0}=-1/4$, but the agreement for $A_{0}=-3/4$ is not quite as good. Further, for the case $A_{0}=-2$ there is no agreement at all, since our asymptotic description (18) predicts that the dispersive waves should not appear for the precise value $A_{0}=-2$, whereas the numerical solutions clearly have waves (albeit of very small amplitude). Of course, since (6) is not of the special class of initial conditions that give rise to $N$-soliton solutions, we know that with (6) there cannot be solutions without dispersive waves. The explanation for the discrepancy is that the exponential asymptotics that led to (18) is only a first approximation. As we vary the parameter $A_{0}$ so that it takes values closer and closer to the triangular numbers N(N+1)-N(N+1), higher-order correction terms will inevitably come into play and eventually dominate.
@@
Indeed, extending the local behaviour (3) of the initial condition about its singularity to be
$u_{0}\sim\frac{A_{0}}{(x-x_{0})^{2}}+\frac{A_{1}}{x-x_{0}}+A_{2}+A_{3}(x-x_{0})+\ldots\quad\mbox{as}\quad x\rightarrow x_{0},$ (21)
@@
the full asymptotics for the dispersive waves will depend not only on $A_{0}$, but also (linearly) on $A_{1}$, $A_{2}$ and $A_{3}$, and so on, in an increasingly complicated manner. For our special initial condition (4), an expansion about $x_{0}=\pi\mathrm{i}/2$ shows that $A_{1}=0$, so the first-order correction terms vanish, while for (6) an expansion about $x_{0}=\mathrm{i}$ shows that $A_{1}=\mathrm{i}A_{0}$, so the first-order correction terms in this case do not vanish. To see the effect of the difference between these two cases, we show in figure (3)(b) a plot of the scaled amplitude versus $A_{0}$ for the initial condition (6). The agreement between the asymptotic prediction (in red) and the numerics (blue dots) appears to be good only for small values of $A_{0}$. Compared with figure (3)(a) (for which $A_{1}=0$), it is clear that higher-order corrections (that depend linearly on $A_{1}$) are significant for (6) (for which $A_{1}=\mathrm{i}A_{0}$) as $|A_{0}|$ increases. To support this conclusion, we test a further initial condition, namely
$u_{0}=-\frac{4A_{0}}{(1+x^{2})^{2}}+\frac{2A_{0}}{1+x^{2}},$ (22)
@@
This is a carefully constructed initial condition that has the same leading-order behaviour near $x_{0}=\mathrm{i}$ as (6), but has an additional term added so that $A_{1}=0$. Thus, locally near $x_{0}=\mathrm{i}$, the initial condition (22) is acting more like the special case (4) near $x_{0}=\pi\mathrm{i}/2$ (which also has $A_{1}=0$). We plot in figure (3)(c) the scaled amplitude versus $A_{0}$ for (22), which shows much better agreement between numerics and asymptotics than in figure 3(b), again reinforcing the conclusion that there are important correction terms that scale with $A_{1}$. While the treatment of correction terms in the context of our exponential asymptotics is rather complicated, we shall summarise how the analysis works for the special case $A_{0}=-N(N+1)$ in appendix A.4.
The plots in figures 24 are generated for a fixed small time. As a simple check of the temporal scaling, we track the first nine dispersive wave crests for a specific example in figure 5. An image of the wave profile is shown in part (a) and the numerically-determined crest location as a function of time is plotted in (b) for the first nine crests. The latter results reformatted on a log-log plot in (c) supports the $t^{1/3}$ scaling. Further, in part (d) we have plotted the asymptotic prediction (19) for the first nine crests $m=1,\ldots,9$. This prediction shows good agreement with the numerics, especially given it holds formally for large $m$ (and small $t$) whereas these numerical results are for small $m$.
@@
Figure 5: (a) A snapshot of a numerical solution of (1) with the initial condition (6), plotted for $t=0.3$ with the first nine wave crests indicated by red dots. (b)–(c) Numerically-determined crest locations as a function of time (solid red) together with the asymptotic prediction (19) (blue dots) for $m=1,\ldots,9$. The slope of the hypotenuse of the triangle in the log-log plot indicates the scaling $|x_{m}|\sim\,\mathrm{constant}\,t^{1/3}$.
@@
For the initial conditions we are considering here, there is a maximum (a “peak”) at $x=0$. This main peak can initially move to either the left or the right, depending on initial condition and the value $A_{0}$. For example, we show in figure 6 the small-time evolution of the main peak for the initial condition (6). Note the behaviour of the solution in the neighbourhood of this peak is not driven by the exponentially-small terms (18), as these are only relevant to the left of $x=-y_{0}/\sqrt{3}=-1/\sqrt{3}$ (since the singularity is at $x=\mathrm{i}$ in this case) in the limit $t\rightarrow 0^{+}$. Instead, for very small times, we expect the peak to move according to the first couple of terms in the power series, namely (7), which suggests the peak location behaves like $x_{\mathrm{peak}}\sim-6(4A_{0}+3)t$. That is, we expect the peak to initially move to the left for $A_{0}<-3/4$ and to the right for $3/4<A_{0}<0$. We can see this behaviour in the left panel of figure 6, where the borderline case $A_{0}=-3/4$ is evident. In the right panel, these numerical results on a log-log plot support the asymptotic prediction by approaching the appropriate straight line with slope unity for large negative values of $\ln t$. Thus, an interesting observation is that the drift of the main peak of the solution scales like $t$ in the small time limit, found by tracking the local maximum of the first couple of terms in the power series expansion (7), while the crests of the dispersive waves scale like $t^{1/3}$ (via (19)), driven by the exponentially-small terms that appear beyond all orders of the power series.
@@
Figure 6: (left) The location of the main peak in real solutions of (1) (horizontal axis) plotted against time (vertical axis). In black, numerical results are shown for the initial condition (6) plotted for $A_{0}=-0.05$, 0.15-0.15, 0.25-0.25, 0.5-0.5, 0.625-0.625, 0.75-0.75, 1-1 and 2-2. (right) Main peak versus time on a log-log plot for the same initial condition, which shows how the numerical results (thick curves) approach the predicted limiting behaviour that comes from analysing the first two terms in the power series (7), namely $\ln t\sim\ln|x_{\mathrm{peak}}|-\ln|6(4A_{0}+3)|$ as $t\rightarrow 0^{+}$ (thin solid lines). Importantly (and as might be expected), the small-time behaviour of the main peak does not come from the exponential terms that appear beyond all order of the power series.
@@
In summary, for our KdV problem (1), we see that the emergence of dispersive waves for small time can be explained by observing how exponentially-small terms are switched on across the point at which the Stokes lines intersect the real-$x$ axis. As far as we are aware, this is the first asymptotic description of how dispersive waves propagate for the KdV equation in the limit that $t\rightarrow 0^{+}$. While these results are interesting in their own right, we also use the above information about the Stokes line structure to inform how we apply far-field conditions in the inner problem in subsection 3.1. Further, our analysis here locates an anti-Stokes lines at $\mathrm{arg}(x-x_{0})=-\pi$, which is (near) where we expect there to be complex-plane singularities propagating out from each singularity of the initial condition, $x=x_{0}$. One of our goals for the remainder of this paper is to track the initial dynamics of the singularities born at $x=x_{0}$, including the important string of singularities that align (close to) $\mathrm{arg}(x-x_{0})=-\pi$.

### v2

@@ 2.2 Exponential asymptotics argument for limit $t\rightarrow 0^{+}$
$\times\cos\left(\frac{2}{3(3t)^{1/2}}(x^{2}+y_{0}^{2})^{3/4}\sin{\textstyle\frac{3}{2}}\phi+{\textstyle\frac{1}{4}}\phi\right),$ (2.12)
@@
these waves take the form (1.5). Thus, we see dispersive waves whose amplitude is exponentially small in both limits $x\rightarrow$-$\infty$ and $t\rightarrow 0^{+}$ (compared to the initial condition $u_{0}(x)$, which is assumed to be $\mathcal{O}(1)$), scaling as
@@
$x_{m}\sim-3\pi^{2/3}m^{2/3}t^{1/3},\quad m\rightarrow\infty,\quad t\rightarrow 0^{+},$ (2.13)
so that the speed of the waves decreases algebraically as $\dot{x}_{m}\sim-\pi^{2/3}(t/m)^{-2/3}$ and the wavelength increases as $x_{m+1}+x_{m}\sim 2\pi^{2/3}(t/m)^{1/3}$ as $t$ increases from zero. These scalings explain why it is so difficult to compute numerical solutions of the KdV model (1.1) accurately for small time using elementary techniques (such as finite differences), given the challenge of capturing fast-moving, small-amplitude waves that would immediately reflect off the boundary of any truncated domain. Such issues can be resolved, for example, via an inverse-scattering formulation, but such an approach is highly nontrivial [92] (and, of course, limited to integrable systems); we instead use the spin (stiff pde integrator) command [75] in Chebfun [40], which is good enough for our real-valued numerical solutions.
@@ 2.3 Numerical tests for the exponential asymptotics
To test (2.12) against numerics, we consider the small-time expansion
@@
i.e., we retain only the first two terms of the algebraic expansion (2.1) but include the leading-order exponential correction (2.12) (in order to capture the dominant oscillatory contribution). For the initial condition (1.4), we have
@@
and $y_{0}=1$. We also make use of a scaled version of (2.12), namely $U_{\mbox{\footnotesize scaled}}=U_{\mbox{\footnotesize dis}}/K$, where
$K=t^{-3/4}(x^{2}+y_{0}^{2})^{1/8}\mathrm{e}^{2(x^{2}+y_{0}^{2})^{3/4}\cos(3\phi/2)/(3(3t)^{1/2})}$ (2.14)
@@
For example, in figure 2(a), we show a numerical solution of (1.1), $u_{\mathrm{num}}$, computed with the $\mathrm{sech}^{2}$-type initial condition (1.4), with $A_{0}=-1/4$, for a representative early time, $t=0.02$. The very small dispersive waves we see in the inset on the left panel are propagating to the left. In the middle panel, we plot $u_{\mathrm{num}}-(u_{0}+tu_{1})$ at the specific time $t=0.02$, and compare with $U_{\mbox{\footnotesize dis}}$ from (2.12). We see for these parameter values, the comparison is very good, which is a strong test for both our numerics and asymptotics. As another test, we plot in the right panel the scaled versions $u_{\mathrm{num}}-(u_{0}+tu_{1})/K$ and $U_{\mbox{\footnotesize scaled}}=U_{\mbox{\footnotesize dis}}/K$, where $K$ is defined in (2.14). The agreement is excellent.
@@
Figure 2: (a) [left panel] Numerical solution of the KdV model (1.1) on the real line, computed at $t=0.02$ using the $\mathrm{sech}^{2}$-type initial condition (1.4) with $A_{0}=-1/4$; [middle panel] plots of $u_{\mathrm{num}}-(u_{0}+tu_{1})$ (blue) and $U_{\mbox{\footnotesize dis}}$ (red circles) versus $x$ using same parameters as in left panel; [right panel] plots of $\left(u_{\mathrm{num}}-(u_{0}+tu_{1})\right)/K$ (blue) and $U_{\mbox{\footnotesize scaled}}$, again with same parameters as left panel. (b) Same as (a), except that $A_{0}=-3/4$.
We have generated a number of other examples like that presented in figure 2(a) using the initial condition (1.4), computed for other values of $A_{0}<0$, and the agreement between numerics and asymptotics is also excellent. One such example is for $A_{0}=-3/4$, is shown in figure 2(b). This sweep of parameters includes the special cases $A_{0}=-N(N+1)$, where $N\in\mathbb{N}$, for which we have exact $N$-soliton solutions with no dispersive waves at all. For those special choices, the amplitude of the waves is predicted by (2.12) to vanish, since
$\cos\left(\frac{\pi}{2}\sqrt{1-4A_{0}}\right)=\cos\left(\frac{\pi}{2}\sqrt{(2N+1)^{2}}\right)=0$
for all $N\in\mathbb{N}$. Putting it together, we have presented in figure 3(a) a comparison of our numerical estimate of the amplitude of the wavelike term $\left(u_{\mathrm{num}}-(u_{0}+tu_{1})\right)/K$ (blue dots) with the asymptotic prediction $(2/3^{3/4}\pi^{1/2})\cos\left(\frac{\pi}{2}\sqrt{1-4A_{0}}\right)$ (red solid) for simulations that take the $\mathrm{sech}^{2}$-type initial condition (1.4). Given the very good agreement in this plot over a range of values of $A_{0}$, we are confident the approximation (2.12) is correctly describing the dispersive waves, at least for the initial condition (1.4), in the small-time limit.
@@
Figure 3: A numerically computed amplitude of the scaled dispersive waves $\left(u_{\mathrm{num}}-(u_{0}+tu_{1})\right)/K$ (blue dots) plotted for various values of the parameter $A_{0}$ at a fixed time $t=0.03$, compared with the asymptotic prediction $(2/3^{3/2}\pi^{1/2})\cos\left(\frac{\pi}{2}\sqrt{1-4A_{0}}\right)$ (red solid). (a) the $\mathrm{sech}^{2}$-type initial condition (1.4); (b) the initial condition (1.6); (c) the refined initial condition (2.16).
@@
Figure 4: (a) [left panel] Numerical solution of the KdV model (1.1) on the real line, computed at $t=0.02$ using the generic initial condition (1.6) with $A_{0}=-1/4$; [middle panel] plots of $u_{\mathrm{num}}-(u_{0}+tu_{1})$ (blue) and $U_{\mbox{\footnotesize dis}}$ (red circles) versus $x$ using same parameters as in left panel; [right panel] plots of $\left(u_{\mathrm{num}}-(u_{0}+tu_{1})\right)/K$ (blue) and $U_{\mbox{\footnotesize scaled}}$, again with same parameters as left panel. (b) Same as (a), except that $A_{0}=-3/4$. (c) Same as (a), except that $A_{0}=-2$.
@@
These types of numerical tests turn out to be slightly more subtle for the initial condition (1.6). For example, in figure 4 we show plots that are analogous to figure 2, except that in figure 4 we use the initial condition (1.6). Broadly speaking, we see very good agreement between numerics and asymptotics for $A_{0}=-1/4$, but the agreement for $A_{0}=-3/4$ is not quite as good. Further, for the case $A_{0}=-2$ there is no agreement at all, since our asymptotic description (2.12) predicts that the dispersive waves should not appear for the precise value $A_{0}=-2$, whereas the numerical solutions clearly have waves (albeit of very small amplitude). Of course, since (1.6) is not of the special class of initial conditions that give rise to $N$-soliton solutions, we know that with (1.6) there cannot be solutions without dispersive waves. The explanation for the discrepancy is that the exponential asymptotics that led to (2.12) is only a first approximation. As we vary the parameter $A_{0}$ so that it takes values closer and closer to the triangular numbers N(N+1)-N(N+1), higher-order correction terms will inevitably come into play and eventually dominate.
@@
Indeed, extending the local behaviour (1.3) of the initial condition about its singularity to be
$u_{0}\sim\frac{A_{0}}{(x-x_{0})^{2}}+\frac{A_{1}}{x-x_{0}}+A_{2}+A_{3}(x-x_{0})+\ldots\quad\mbox{as}\quad x\rightarrow x_{0},$ (2.15)
@@
the full asymptotics for the dispersive waves will depend not only on $A_{0}$, but also (linearly) on $A_{1}$, $A_{2}$ and $A_{3}$, and so on, in an increasingly complicated manner. For our special initial condition (1.4), an expansion about $x_{0}=\pi\mathrm{i}/2$ shows that $A_{1}=0$, so the first-order correction terms vanish, while for (1.6) an expansion about $x_{0}=\mathrm{i}$ shows that $A_{1}=\mathrm{i}A_{0}$, so the first-order correction terms in this case do not vanish. To see the effect of the difference between these two cases, we show in figure (3)(b) a plot of the scaled amplitude versus $A_{0}$ for the initial condition (1.6). The agreement between the asymptotic prediction (in red) and the numerics (blue dots) appears to be good only for small values of $|A_{0}|$. Compared with figure (3)(a) (for which $A_{1}=0$), it is clear that higher-order corrections (that depend linearly on $A_{1}$) are significant for (1.6) (for which $A_{1}=\mathrm{i}A_{0}$) as $|A_{0}|$ increases. To support this conclusion, we test a further initial condition, namely
$u_{0}=-\frac{4A_{0}}{(1+x^{2})^{2}}+\frac{2A_{0}}{1+x^{2}},$ (2.16)
@@
This is a carefully constructed initial condition that has the same leading-order behaviour near $x_{0}=\mathrm{i}$ as (1.6), but has an additional term added so that $A_{1}=0$. Thus, locally near $x_{0}=\mathrm{i}$, the initial condition (2.16) is acting more like the special case (1.4) near $x_{0}=\pi\mathrm{i}/2$ (which also has $A_{1}=0$). We plot in figure (3)(c) the scaled amplitude versus $A_{0}$ for (2.16), which shows much better agreement between numerics and asymptotics than in figure 3(b), again reinforcing the conclusion that there are important correction terms that scale with $A_{1}$. While the treatment of correction terms in the context of our exponential asymptotics is rather complicated, we shall summarise how the analysis works for the special case $A_{0}=-N(N+1)$ in appendix A.4.
The plots in figures 24 are generated for a fixed small time. As a simple check of the temporal scaling, we track the first nine dispersive wave crests for a specific example in figure 5. An image of the wave profile is shown in part (a) and the numerically-determined crest location as a function of time is plotted in (b) for the first nine crests. The latter results reformatted on a log-log plot in (c) supports the $t^{1/3}$ scaling. Further, in part (d) we have plotted the asymptotic prediction (2.13) for the first nine crests $m=1,\ldots,9$. This prediction shows good agreement with the numerics, especially given it holds formally for large $m$ (and small $t$) whereas these numerical results are for small $m$.
@@
Figure 5: (a) A snapshot of a numerical solution of (1.1) with the initial condition (1.6), plotted for $t=0.3$ with the first nine wave crests indicated by red dots. (b)–(c) Numerically-determined crest locations as a function of time (solid red) together with the asymptotic prediction (2.13) (blue dots) for $m=1,\ldots,9$. The slope of the hypotenuse of the triangle in the log-log plot indicates the scaling $|x_{m}|\sim\,\mathrm{constant}\,t^{1/3}$.
@@
For the initial conditions we are considering here, there is a maximum (a “peak”) at $x=0$. This main peak can initially move to either the left or the right, depending on initial condition and the value $A_{0}$. For example, we show in figure 6 the small-time evolution of the main peak for the initial condition (1.6). Note the behaviour of the solution in the neighbourhood of this peak is not driven by the exponentially-small terms (2.12), as these are only relevant to the left of $x=-y_{0}/\sqrt{3}=-1/\sqrt{3}$ (since the singularity is at $x=\mathrm{i}$ in this case) in the limit $t\rightarrow 0^{+}$. Instead, for very small times, we expect the peak to move according to the first couple of terms in the power series, namely (2.1), which suggests the peak location behaves like $x_{\mathrm{peak}}\sim-6(4A_{0}+3)t$. That is, we expect the peak to initially move to the left for $A_{0}<-3/4$ and to the right for $3/4<A_{0}<0$. We can see this behaviour in the left panel of figure 6, where the borderline case $A_{0}=-3/4$ is evident. In the right panel, these numerical results on a log-log plot support the asymptotic prediction by approaching the appropriate straight line with slope unity for large negative values of $\ln t$. Thus, an interesting observation is that the drift of the main peak of the solution scales like $t$ in the small time limit, found by tracking the local maximum of the first couple of terms in the power series expansion (2.1), while the crests of the dispersive waves scale like $t^{1/3}$ (via (2.13)), driven by the exponentially-small terms that appear beyond all orders of the power series.
@@
Figure 6: (left) The location of the main peak in real solutions of (1.1) (horizontal axis) plotted against time (vertical axis). In black, numerical results are shown for the initial condition (1.6) plotted for $A_{0}=-0.05$, 0.15-0.15, 0.25-0.25, 0.5-0.5, 0.625-0.625, 0.75-0.75, 1-1 and 2-2. (right) Main peak versus time on a log-log plot for the same initial condition, which shows how the numerical results (thick curves) approach the predicted limiting behaviour that comes from analysing the first two terms in the power series (2.1), namely $\ln t\sim\ln|x_{\mathrm{peak}}|-\ln|6(4A_{0}+3)|$ as $t\rightarrow 0^{+}$ (thin solid lines). Importantly (and as might be expected), the small-time behaviour of the main peak does not come from the exponential terms that appear beyond all order of the power series.
@@
In summary, for our KdV problem (1.1), we see that the emergence of dispersive waves for small time can be explained by observing how exponentially-small terms are switched on across the point at which the Stokes lines intersect the real-$x$ axis. As far as we are aware, this is the first asymptotic description of how dispersive waves propagate for the KdV equation in the limit that $t\rightarrow 0^{+}$. While these results are interesting in their own right, we also use the above information about the Stokes line structure to inform how we apply far-field conditions in the inner problem in subsection 3.1. Further, our analysis here locates an anti-Stokes lines at $\mathrm{arg}(x-x_{0})=-\pi$, which is (near) where we expect there to be complex-plane singularities propagating out from each singularity of the initial condition, $x=x_{0}$. One of our goals for the remainder of this paper is to track the initial dynamics of the singularities born at $x=x_{0}$, including the important string of singularities that align (close to) $\mathrm{arg}(x-x_{0})=-\pi$.